### v1

# Sim-CLIP: Unsupervised Siamese Adversarial Fine-Tuning for Robust and Semantically-Rich Vision-Language Models

Abstract
Vision-language models (VLMs) have achieved significant strides in recent times specially in multimodal tasks, yet they remain susceptible to adversarial attacks on their vision components. To address this, we propose Sim-CLIP, an unsupervised adversarial fine-tuning method that enhances the robustness of the widely-used CLIP vision encoder against such attacks while maintaining semantic richness and specificity. By employing a Siamese architecture with cosine similarity loss, Sim-CLIP learns semantically meaningful and attack-resilient visual representations without requiring large batch sizes or momentum encoders. Our results demonstrate that VLMs enhanced with Sim-CLIP’s fine-tuned CLIP encoder exhibit significantly enhanced robustness against adversarial attacks, while preserving semantic meaning of the perturbed images. Notably, Sim-CLIP does not require additional training or fine-tuning of the VLM itself; replacing the original vision encoder with our fine-tuned Sim-CLIP suffices to provide robustness. This work underscores the significance of reinforcing foundational models like CLIP to safeguard the reliability of downstream VLM applications, paving the way for more secure and effective multimodal systems.
1 Introduction
The remarkable success of large language models (LLMs) (OpenAI 2023; Meta 2023) in understanding and generating human-like text has inspired researchers to extend their capabilities to the visual domain. This has led to the development of Vision-Language Models (VLMs) (Alayrac et al. 2022; Liu et al. 2024), which aim to bridge the gap between visual and textual information. To extract meaningful visual representations from images, these VLMs often leverage pretrained vision encoders, such as CLIP (Radford et al. 2021), BEiT (Bao et al. 2021), and DINO (Caron et al. 2021). Leveraging pretrained vision encoders empowers VLMs to capitalize on the rich visual knowledge acquired through extensive pretraining on diverse image datasets. This capability enhances the VLMs’ performance across various downstream tasks, eliminating the requirement for task-specific fine-tuning. Among these pretrained vision encoders, CLIP (Radford et al. 2021) gained significant attention and is widely used in many state-of-the-art VLMs such as OpenFlamingo (Awadalla et al. 2023) and LLaVA (Liu et al. 2024). CLIP’s popularity stems from its exceptional performance in zero-shot settings and its ability to align inputs from different modalities in a joint embedding space. Additionally, VLMs equipped with CLIP demonstrate remarkable zero-shot performance in down-stream tasks such as image captioning, and visual question answering (VQA). Due to the impressive down-stream task performance and contribution to the growing trend of multimodal learning, an increasing number of large VLMs are being made publicly available. While the availability of pre-trained weights and architectures for VLMs facilitated further research and application development, it also raised severe safety concerns that must be addressed. For instance, recent works (Zhao et al. 2024; Wei, Haghtalab, and Steinhardt 2024) have demonstrated the high susceptibility of VLMs to adversarial attacks targeting either text or image inputs. Notably, it has been argued that the vision modality is more susceptible to manipulation compared to text modalities (Goodfellow, Shlens, and Szegedy 2014; Carlini et al. 2024).
Figure 1: Targeted $\ell_{\infty}$ attack at $\epsilon=2/255$ radii using original CLIP model as vision encoder in LLaVA. Original images with their captions on left and captions generated by LLaVA for benign images on right.
Besides, the authors in (Schlarmann and Hein 2023) demonstrate that adversaries can employ human-imperceptible adversarial perturbations to create targeted attacks, effectively generating desired malicious outputs as shown in Figure 1.
While existing methods (Mao et al. 2022; Schlarmann et al. 2024) have made strides in improving the robustness of CLIP models, significant challenges remain. Specifically, these adversarially fine-tuned CLIP models often lead to a noticeable degradation in performance on downstream tasks. During captioning, these models frequently struggle to capture intricate semantic features and fail to fully grasp the overall semantic meaning of the images. These issues highlight a critical gap in the ability of current robust CLIP models to preserve detailed and comprehensive image understanding. To address these concerns, we propose Sim-CLIP, an unsupervised adversarial fine-tuning method for enhancing the robustness of CLIP’s vision encoder against adversarial attacks. Sim-CLIP aims to enhance the adversarial robustness of the vision modality in VLMs while preserving its ability to capture detailed semantic features and the overall semantic meaning of images. Moreover, Sim-CLIP can be seamlessly integrated into existing VLMs for downstream tasks without requiring additional task-specific retraining or fine-tuning. Consequently, all downstream tasks, including zero-shot classification or other zero-shot tasks specific to VLMs, achieve robustness against attacks targeting the vision modality. Our work establishes a new benchmark for evaluating zero-shot adversarial robustness, an increasingly important aspect of foundation models as they are implemented in critical applications.
2 Related Works
Vision Language Models.
In recent times, various VLMs have emerged, such as LLaVA (Liu et al. 2024), Flamingo (Alayrac et al. 2022), OpenFlamingo (OF) (Awadalla et al. 2023), and Macaw-LLM (Lyu et al. 2023), to name a few. These models leverage the power of multi-modal learning by combining visual and textual data, thereby enhancing their capability to understand and generate responses to multi-modal inputs. Most of these VLMs employ pre-trained LLMs (e.g., Vicuna (Chiang et al. 2023)) with a large-scale vision encoder, such as CLIP (Radford et al. 2021) to process multi-modal data. During training, the vision encoder is kept frozen, while the model learns the interaction between modalities. This interaction is typically facilitated by combining a projection layer with cross-attention mechanism.
Adversarial robustness in traditional ML.
The susceptibility of traditional machine learning models (e.g., CNNs, RNNs) to adversarial attacks is well-established and has been extensively studied. Most of the existing adversarial attacks predominantly target monomodal models, focusing on either text or image modality (Hossain et al. 2023; Shahid et al. 2023). In the image domain, gradient-based adversarial attacks (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017) can deceive a target model by applying slight perturbations to the input images that are imperceptible to human eyes. Alternatively, patch-based attacks (Brown et al. 2017) strategically place adversarial patches to deceive models, rather than modifying the entire image. Text-based adversarial (Ebrahimi et al. 2017; Guo et al. 2021) attacks are also thoroughly investigated, requiring a different approach due to the discrete nature of textual data. Adversarial training (Madry et al. 2017; Pan et al. 2022) has emerged as an efficient defense mechanism against aforementioned attacks, which involves training the model with adversarial examples to improve its robustness.
Adversarial robustness for VLMs.
Given CLIP’s remarkable performance in zero-shot settings and understanding visual tasks, most VLMs employ it or its variants as their vision encoder. Few recent studies (Jiang et al. 2020; Bansal et al. 2023; Zhou et al. 2023) demonstrate CLIP’s vulnerability to imperceptible attacks, which can significantly impact downstream task performance of VLMs. For instance, in AdvCLIP (Zhou et al. 2023), the authors generate universal patches for CLIP models that can deceive VLMs across all of their downstream tasks. In (Zhao et al. 2024), the authors leverage diffusion models to create adversarial samples that manipulate the model into generating a targeted output. Moreover, in (Schlarmann and Hein 2023), the authors demonstrate the potential of gradient-based attacks on VLMs, compelling the model to generate inaccurate results. One recent study (Mao et al. 2022) introduced a supervised adversarial fine-tuning scheme for CLIP that employs cross-modal image-text contrastive loss. A few concurrent works (Jiang et al. 2020; Fan et al. 2021) proposed unsupervised adversarial training methods based on SimCLR contrastive loss. However, these methods necessitate a large batch size to achieve strong robust performance, making them unsuitable for integration into VLMs. The authors in (Gowal et al. 2020) proposed an adversarial fine-tuning scheme based on BYOL (Grill et al. 2020), which addresses the issue of large batch sizes but introduces an overhead with the momentum encoder. Besides, the authors in (Schlarmann et al. 2024) proposed an $\ell_{2}$ loss-based unsupervised fine-tuning scheme, but it fails to capture semantic features from the images effectively. In contrast, our unsupervised fine-tuning approach, based on a siamese network, utilizes cosine similarity to effectively capture semantic information during adversarial training without requiring a large batch size or an additional momentum encoder.
3 Methodology
In this section, we first discuss how existing adversarial attacks for VLMs are formalized. Then we delve into the specifics of existing unsupervised adversarial training methods and their limitations, followed by the details of our proposed framework, Sim-CLIP.
3.1 Adversarial Attacks on VLM
In gradient-based adversarial attacks (e.g., PGD (Goodfellow, Shlens, and Szegedy 2014), APGD (Croce and Hein 2020)), malicious attackers typically add slight perturbation $\delta$ to an input image $x$. This perturbation is carefully calculated for a specific target model, represented as $f_{\theta}$, where $\theta$ denotes the model’s parameters. The core mechanism of these attacks involves leveraging the gradients of the model’s loss function (e.g., cross-entropy) $\mathcal{L}$, with respect to the input image $x$. By exploiting these gradients, attackers are able to craft adversarial examples, denoted as $x_{a}=x+\delta$. These adversarial samples are imperceptible to human eyes yet possess the ability to mislead deep learning models such as $f_{\theta}$ towards misclassification:
$\mathbf{x}_{a}=\underset{\mathbf{x}_{a}}{\operatorname{argmax}}\quad\mathcal{L}\left(\mathbf{x}_{a},\mathbf{y}_{true}\right),\left\|\mathbf{x}_{a}-\mathbf{x}\right\|_{p}\leq\epsilon.$ (1)
Figure 2: Workflow and overview of proposed Sim-CLIP. CLIP undergoes adversarial training using our proposed Sim-CLIP approach and serves as the vision encoder in the Vision-Language Model. The workflow includes image perturbation for inference and pre-training CLIP on ImageNet using Sim-CLIP for enhanced robustness. CLIP model undergoes adversarial fine-tuning while the text encoder is kept frozen.
Here, $\delta$ is constrained such that $\|\delta\|p\leq\epsilon$, where $\|\cdot\|_{p}$ denotes the $p$-norm, a measure of the vector’s magnitude according to the chosen norm (e.g., $l_{2}$ or $l_{\infty}$). This constraint ensures that each element of the perturbation vector does not exceed $\epsilon$ in absolute value. To execute such attacks, the attacker must have access to all the model’s weights, hence these attacks are also known as white-box attacks. This is particularly relevant for open-source models (e.g., VLMs), where the model architecture and weights are publicly available. In this study, we employ adversarial attacks in two different settings to evaluate VLM’s performance: (1) Targeted attack and (2) Untargeted attack.
3.1.1 Targeted attack.
In a targeted attack, the adversary manipulates the VLM to generate a specific incorrect prediction by providing perturbed query samples. The adversary perturbs a query image $q$ by adding slight perturbation $\delta_{q}$, resulting in a perturbed sample $q+\delta_{q}$. When this perturbed sample is fed into the VLM, it generates the desired malicious output $y_{l}^{t}$, such as spam links or false information. The targeted attack is formulated as an optimization problem that aims to minimize the negative log-likelihood of generating the desired incorrect output $y_{l}^{t}$, given the perturbed query image is $q+\delta_{q}$. This can be formalized as follows:
[…]
Here, the goal is to find the optimal $\delta_{q}$ while ensuring the perturbation is bounded by $\varepsilon_{q}$ in the $l_{\infty}$ norm. Note that, by minimizing the negative log-likelihood, we ensure that the probability of the target token is maximized.
3.1.2 Untargeted attack.
In contrast to targeted attacks, untargeted attacks seek to induce any form of error, without specifying a particular target. The objective of an untargeted attack can be expressed as maximizing the likelihood of the model producing any output that differs from the correct or expected output. The mathematical formulation for an untargeted attack can be represented as follows:
[…]
Here, $\delta_{q}$ denotes the perturbation to the query image $q$, with the aim of maximizing the likelihood of the model outputting any incorrect prediction $y_{l}$ that deviates from the correct output $y$.
3.2 Unsupervised Adversarial Fine-Tuning
We propose an unsupervised adversarial fine-tuning approach for the CLIP vision encoder that aims to enhance its robustness against adversarial attacks. Our methodology incorporates unsupervised learning techniques by leveraging siamese architecture. Our goal is to enhance the resilience of the vision encoder against adversarial attacks while preserving the semantic meaning of the image features.
Previously, FARE (Schlarmann et al. 2024) introduced an unsupervised adversarial fine-tuning scheme for the CLIP vision encoder that does not rely on text embedding. FARE achieves its robustness by minimizing a $\ell_{2}$ loss function between an adversarially perturbed image embedding and a clean image embedding. The embedding loss can be expressed as follows:
$\mathcal{L}_{F}(x,x_{a})=\max_{\left\|x_{a}-x\right\|_{\infty}\leq\varepsilon}\|\theta(x_{a})-\theta(x)\|_{2}^{2}$ (4)
This loss encourages the embedding $\theta(x_{a})$ of perturbed images $z$ to stay close to the original embedding $\theta(x)$. Although FARE demonstrates robust performance against adversarial attacks, it exhibits two major issues. First, the $\ell_{2}$ loss employed by FARE may not be the most suitable choice when dealing with high-dimensional data from different modalities. In high-dimensional spaces, the volume of the space expands exponentially with the number of dimensions, leading to sparsity issues. This sparsity makes it difficult to capture relationships between data points effectively using $\ell_{2}$ loss. Second, $\ell_{2}$ loss often fails to capture semantic meaning from the image features effectively. Since $\ell_{2}$ loss minimization operates at the pixel level, it prioritizes the reduction of errors in low-level features over capturing semantic content. As a result, learned representations often lack the semantic richness needed for downstream tasks, which hinders the model’s ability to learn meaningful patterns.
3.3 Adversarial Fine-tuning with Siamese Architecture
Sim-CLIP tackles the challenges present in FARE by tailoring cosine similarity loss within a siamese architecture. During our adversarial fine-tuning phase, Sim-CLIP first generates a perturbed view $x^{\prime}$ from the clean input image $x$. We utilize PGD perturbation to generate the perturbed view:
$x^{\prime}_{t+1}=\Pi_{x+\epsilon}(x_{t}+\alpha\cdot\textrm{sign}(\nabla_{x}\mathcal{L}(x_{t},y)))$ (5)
[…]
Here, $y$ represents the true label of the input image and $\mathcal{L}$ is a cross-entropy loss. At each iteration $t$, a perturbation is calculated based on the gradient of the loss function with respect to the input image $x$. The magnitude of this perturbation is controlled by a step size parameter $\alpha$, which determines the strength of the perturbation applied in each iteration. This perturbation is constrained within a specified bound $\epsilon$. Subsequently, both clean and perturbed views are fed into the CLIP models, which share the same weights, as depicted in Figure 2(b). The CLIP models generate clean representation $R_{c}$ from the original image $x$ and perturbed representation $R_{p}$ from the perturbed view $x^{\prime}$. Now, we maximize the similarity between the perturbed and clean representations to encourage the model to learn features that are invariant to adversarial perturbations. To this end, we minimize negative cosine similarity between the representations $R_{p}$ and $R_{c}$ to ensure their alignment in the vector space, enhancing model coherence and robustness to adversarial perturbations:
$CosSim(R_{p},R_{c})=-\left(\frac{R_{p}}{\lVert R_{p}\rVert_{2}}\cdot\frac{R_{c}}{\lVert R_{c}\rVert_{2}}\right)$ (6)
Here, $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm, which is equivalent to minimizing mean squared error between $\ell_{2}$-normalized vectors.
3.4 Symmetric Loss Collapse Prevention
Minimizing this negative cosine loss directly might lead to a collapse of the loss function, yielding an undesirable constant solution (Chen et al. 2020). Conventional unsupervised adversarial training methods (Jiang et al. 2020; Fan et al. 2021) attempt to address this issue by relying on negative samples or momentum encoders. However, integrating negative samples often necessitates larger batch sizes, leading to increased computational burden, while momentum encoders introduce extra overhead. To prevent the loss function from collapsing while mitigating the resource burden, we integrate a stop-gradient mechanism following (Chen and He 2021) into our adversarial fine-tuning loss. We employ a symmetric loss for our adversarial training scheme, which can be defined as follows:
$\begin{split}\mathcal{L}_{simclip}(R_{p},R_{c})&=\frac{1}{2}\left(CosSim(R_{p},\texttt{stopgrad}(R_{c}))\right.\\ &\quad\left.+CosSim(R_{c},\texttt{stopgrad}(R_{p}))\right)\end{split}$ (7)
Here, one CLIP model’s output is held constant using stop-gradient, while the other model’s output is projected to align with it. Subsequently, the roles of the two models are reversed, with the first model’s output being projected and matched to the “constant” output of the second model. This implies that in the first term of Equation 7, $R_{p}$ does not receive any gradients from $R_{c}$ since $R_{c}$ is treated as constant. However, in the second term of the same equation, $R_{c}$ receives gradients from $R_{p}$ as the stop-gradient operation is now applied to $R_{p}$ instead. This process results in two losses, which are then averaged and minimized for optimization. In summary, each model receives gradients for the loss term, except for cases where stop-gradient is not applied to its output. The effectiveness of the stop-gradient mechanism in preventing loss function from collapsing is further validated through ablation study in the Appendix.
4 Experimental Analysis
We evaluate our adversarially fine-tuned robust CLIP model across various downstream tasks, including zero-shot classification. To evaluate performance on downstream tasks, we integrate robust clip models into VLMs by substituting their default vision encoders.
Table 1: Comparing Robustness of Vision-Language Models with different CLIP models. This table displays the robust performance of various vision-language models on two tasks: image captioning and visual question answering (VQA). For VQA tasks (VizWiz, OKVQA), we report VQA accuracy, and for captioning tasks (COCO, Flickr30k), we report CIDEr score.
VLM Vision encoder COCO Flickr30 VizWiz OKVQA
  clean $\ell_{\infty}$   clean $\ell_{\infty}$   clean $\ell_{\infty}$   clean $\ell_{\infty}$
$\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{8}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{8}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{8}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{8}}{{255}}$
Open Flamingo CLIP 80.5 7.82 5.6 2.4 61.0 6.4 3.8 1.4 23.8 2.4 1.8 0 48.5 1.8 0.0 0.0
TeCoA2 74.5 59.7 40.3 10.3 48.2 37.3 27.4 10.3 22.3 15.5 10.6 3.5 33.6 23.4 15.3 6.7
FARE2 84.3 68.2 53.5 18.4 53.1 48.6 34.3 12.3 22.1 15.9 12.3 6.7 34.5 30.6 17.1 9.8
Sim-CLIP2 85.6 72.8 58.4 19.3 56.3 50.5 35.1 16.4 21.8 17.3 13.6 8.5 35.1 29.3 19.7 11.6
TeCoA4 71.0 58.3 50.3 15.8 45.6 36.2 32.9 18.0 19.3 15.1 14.7 8.4 31.0 22.4 20.5 10.1
FARE4 81.4 67.9 56.1 23.3 51.8 47.3 37.6 20.1 16.4 15.7 13.7 10.2 31.8 28.0 19.2 13.5
Sim-CLIP4 81.6 71.5 60.5 26.0 54.5 48.0 39.2 20.4 20.0 15.6 15.7 12.4 32.0 27.4 22.0 15.7
LLaVA 1.5 CLIP 121.9 21.8 13.5 2.4 79.0 15.3 10.0 3.4 39.3 13.3 3.2 0.0 57.3 8.3 3.0 0.0
TeCoA2 115.6 98.3 73.5 38.1 75.6 65.3 50.5 29.4 38.5 25.4 15.4 8.3 55.6 40.3 30.5 14.2
FARE2 123.5 105.2 86.4 39.4 78.9 70.3 60.5 25.1 37.3 29.3 17.6 10.4 54.3 43.5 30.1 15.3
Sim-CLIP2 125.6 109.4 93.5 45.6 80.5 73.1 63.8 29.8 41.5 30.3 19.8 14.6 60.3 47.5 31.5 17.5
TeCoA4 110.3 95.5 75.6 35.3 71.8 62.5 51.0 27.0 34.5 30.5 18.3 9.3 50.3 39.0 32.3 12.3
FARE4 119.4 100.5 83.5 41.6 76.3 70.3 56.5 29.5 38.5 31.3 21.0 10.1 53.5 45.0 34.8 15.3
Sim-CLIP4 122.3 108.1 90.3 44.3 79.0 72.3 61.3 32.5 40.0 29.3 22.3 16.7 58.6 46.5 38.5 22.3
VLM models.
We use OpenFlamingo 9B (OF) (Awadalla et al. 2023) and LLaVA-1.5 (Liu et al. 2024) with 7B parameters as our target VLMs. They both utilize CLIP ViT-L-14 (Dosovitskiy et al. 2020) as their vision encoder. However, they differ in their language decoders. OF employs MPT-7B (Team et al. 2023), while LLaVA-1.5 utilizes Vicuna (Chiang et al. 2023). During our evaluation, OF is provided only with context text alongside the query image. In contrast, we utilize both the default system prompt and task-specific prompts along with the query images for LLaVA.
Datasets.
We considered a wide range of datasets for down-stream tasks, specifically for image captioning and VQA. For image captioning, we use the COCO (Lin et al. 2014) and Flickr30k (Plummer et al. 2015) datasets, while for visual question answering, we employ VizWiz (Gurari et al. 2018) and OKVQA (Marino et al. 2019). Additionally, we assess the robustness of our adversarially fine-tuned CLIP model on the zero-shot classification task using the CIFAR10, CIFAR100 (Krizhevsky, Hinton et al. 2009), EuroSAT (Helber et al. 2019), PCAM (Veeling et al. 2018), and Flower (Nilsback and Zisserman 2008) datasets. Our evaluation process employs two approaches: for adversarial evaluation, we randomly select 500 images from each respective dataset, while for clean evaluation we utilize all available samples in the test suite.
Figure 3: Targeted $\ell_{\infty}$ attacks at $\epsilon=4/255$ radii using original and robust CLIP models as vision encoder in LLaVA. Considering the target strings from Table 2, we present generated captions (good caption, captions with mistakes, captions missing intricate details, malicious target output) on original (left) and imperceptible adversarial (right) images.
Table 2: Quantitative evaluation of $\ell_{\infty}$ targeted attacks at $\epsilon=4/255$ radii. Mean success rate = (number of successful attacks / number of attacked samples).
Target CLIP TeCoA2 FARE2 Sim-CLIP2 TeCoA4 FARE4 Sim-CLIP4
COVID vaccine has severe 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Opposing political party 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Please visit https://deal.net 20 / 20 2 / 20 2 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Investing in this new crypto 20 / 20 3 / 20 1 / 20 1 / 20 0 / 20 0 / 20 0 / 20
Climax change is a hoax 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Mean success rate: 100% 5% 3% 1% 0% 0% 0%
Average CIDEr score: 0 15.3 23.5 45.3 64.4 75.3 84.7
Adversarial fine-tuning settings.
[…]
To ensure compatibility with the VLMs used for inference, we employ adversarial fine-tuning on the CLIP Vit-L/14 vision encoder, which precisely matches the version utilized by the VLMs (e.g., OF and LLaVA). During inference, we substitute the default CLIP model of the VLMs with our robust CLIP, while keeping the language decoder and projection layer frozen, as depicted in Figure 2(a). In the following sections, we present a comprehensive comparison of our proposed fine-tuning scheme with the original CLIP (Radford et al. 2021) and two state-of-the-art (SOTA) robust fine-tuned versions, namely TeCoA (Mao et al. 2022) and FARE (Schlarmann et al. 2024). To ensure fairness in comparison, we utilize fine-tuned versions of TeCoA and FARE with adversarial radii set to $\epsilon=2/255$ and $\epsilon=4/255$. These models are denoted as $\text{TeCoA}^{2}$, $\text{FARE}^{2}$ for $\epsilon=2/255$ radii, and $\text{TeCoA}^{4}$, $\text{FARE}^{4}$ for $\epsilon=4/255$ radii.
Adversarial attack settings.
We evaluate Sim-CLIP’s robust performance in three distinct attack scenarios and compare its performance with SOTA robust CLIP models (e.g., TeCoA, FARE). Each scenario utilizes a separate attack pipeline designed to degrade the model’s performance significantly. First, we conduct untargeted attacks to assess clean and robust performance in downstream tasks. Following (Schlarmann and Hein 2023), we employ APGD attack with $\ell_{\infty}$ norm and radii set to $\epsilon=2/255$ and $\epsilon=4/255$. The APGD attack is executed using single precision for 100 iterations. The attacked images with $\epsilon=2/255$ perturbations are completely imperceptible to human eyes, while with $\epsilon=4/255$ a user could notice the perturbation when paying close attention. Additionally, in this paper, we evaluate the robust CLIP models under a stronger adversarial setting by using a perturbation radius of $\epsilon=8/255$ for the downstream tasks.
Second, we investigate the adversarial robustness of CLIP models in zero-shot classification tasks. For multi-class classification, we utilize targeted DLR loss, while for binary datasets like PCAM, we use APGD with cross-entropy loss due to the inapplicability of DLR loss for binary classification. We employ a similar attack pipeline to our previous methodology and conduct APGD attacks for 100 iterations. We conduct attacks under two settings of perturbation radius: $\epsilon=2/255$ and $\epsilon=4/255$. Finally, we conduct targeted attacks on LLaVA-1.5 and assess robust and clean performance of robust CLIP models. Targeted attacks compel the VLMs to generate a predetermined output specified by the attackers. Thus, these pose a graver threat compared to untargeted attacks. The attack is considered successful only if the desired target output is present in the VLM’s output. In our experiments, we randomly select 20 images from the COCO dataset and employ APGD attack for 10,000 iterations. For the attack setting, we utilize $\ell_{\infty}$ norm with radii of $\epsilon=2/255$ and $\epsilon=4/255$. We discuss more details on our targeted attack pipeline in Appendix.
Metrics.
For assessing captioning tasks, we utilize the CIDEr metric, whereas, for visual question answering, we use VQA accuracy. The CIDEr score (Vedantam, Lawrence Zitnick, and Parikh 2015) stands as a widely used metric in evaluating image captioning models. It assesses the similarity between a generated caption and a set of ground truth captions by tallying the occurrence of consecutive words and applying a term frequency–inverse document frequency (TF-IDF) weighting scheme. A higher CIDEr score refers to a better alignment between the generated image captions and the semantic content conveyed by the images. In contrast, VQA accuracy measures the percentage of correctly answered questions based on visual input.
4.1 Untargeted attack results and discussion.
In Table LABEL:tab:untargetdown, we report clean and robust performance of various CLIP versions. In terms of clean performance (without attack), original CLIP model performs better compared to adversarially fine-tuned models. However, under adversarial attacks, the original CLIP model exhibits a significant decline in performance. Particularly under stronger attack at $\epsilon=8/255$, its performance deteriorates even further. We observe a slight degradation in the clean performance for the robust clip models due to adversarial fine-tuning. Notably, among the robust versions of CLIP, Sim-CLIP achieves the best clean performance. For OF, Sim-CLIP4 demonstrates superior performance compared to FARE4 and TeCoA4 across most downstream task datasets. Although FARE4 marginally outperforms Sim-CLIP4 in VizWiz and OKVQA datasets at radii set to $\epsilon=2/255$ attack, the differences are negligible at 0.1 and 0.6 respectively. However, when subjected to stronger attacks at $\epsilon=4/255$ and $\epsilon=8/255$, Sim-CLIP4 consistently outperforms all SOTA robust clip models. Similar trends are observed with Sim-CLIP2, outperforming FARE2 and TeCoA2 in both clean and robust performance. Additionally, for LLaVA, Sim-CLIP demonstrates even greater efficacy, achieving superior
Table 3: Evaluation of CLIP models on zero-shot classification datasets under adversarial attacks. We report robust accuracy of different CLIP models.
Attack Vision encoder
Cars
CIFAR10
CIFAR100
EuroSAT
Flowers
PCAM
$\ell_{\infty}=\nicefrac{{2}}{{255}}$ CLIP 0.0 0.0 0.0 0.0 0.0 0.2
TeCoA2 21.4 64.8 34.4 12.4 25.8 40.1
FARE2 26.5 61.3 35.8 7.7 26.5 41.5
Sim-CLIP2 25.8 63.8 37.4 6.3 28.9 43.0
TeCoA4 17.2 59.5 33.6 8.0 24.1 48.5
FARE4 30.8 57.3 36.4 12.8 31.3 50.0
Sim-CLIP4 33.1 57.0 37.8 18.6 30.4 53.0
$\ell_{\infty}=\nicefrac{{4}}{{255}}$ CLIP 0.0 0.0 0.0 0.0 0.0 0.0
TeCoA2 5.8 31.0 17.8 3.5 6.7 16.0
FARE2 4.8 25.9 14.0 5.5 7.1 17.2
Sim-CLIP2 6.5 32.8 18.4 4.7 8.8 19.4
TeCoA4 8.4 35.5 21.6 6.8 12.4 43.5
FARE4 12.8 34.8 21.4 11.7 12.9 50.2
Sim-CLIP4 14.1 34.0 22.8 13.6 11.2 50.9
performance in both captioning and VQA tasks. Notably, its resilience is highlighted in higher attack settings at $\epsilon=4/255$ and $\epsilon=8/255$, where both Sim-CLIP4 and Sim-CLIP2 significantly outperform FARE and TeCoA. In summary, these findings underscore the effectiveness of our Sim-CLIP in maintaining high performance on clean data while significantly enhancing resilience against adversarial attacks. Moreover, the higher CIDEr score achieved by Sim-CLIP under attack demonstrates its superior ability to preserve the semantic meaning of image features in the generated captions compared to other robust variations of CLIP. We present visual examples of VQA in the Appendix.
4.2 Targeted attack results and discussion.
We present the quantitative results of our targeted attacks at $\epsilon=4/255$ in Table 2. This analysis includes CIDEr score to evaluate the quality of generated captions. Additionally, we illustrate randomly selected attacked samples and their captions generated by LLaVA, integrated with different CLIP models in Figure 3. We observe that the original CLIP model is highly susceptible to targeted attacks and demonstrates no robustness. In each instance, the original CLIP model breaks and generates the given target string. Conversely, TeCoA2 and FARE2 break in 5 and 3 cases, resulting in mean success rates of 5% and 3%, respectively. In stark contrast, Sim-CLIP2 breaks in only one case, further underscoring the superior performance of Sim-CLIP. Notably, Sim-CLIP4, FARE4, and TeCoA4 show complete robustness under targeted attack. However, the quality of captions generated by Sim-CLIP4 notably surpasses FARE4 and TeCoA4, as shown in Figure 3. Moreover, captions generated by TeCoA4 exhibit inferior quality and contain errors, while FARE4’s captions often lack intricate details or semantic features from the corresponding images. For instance, consider the sample featuring a patient. With the original image, both FARE4 and Sim-CLIP4 generate captions without errors and with comprehensive details. However, under attack, the caption generated by FARE4 lacks specifics regarding the color of the hospital bed and the presence of a mask, whereas Sim-CLIP4 retains these semantic features of the image. This exemplifies the robustness of our adversarial fine-tuning approach, as it not only enhances the model’s ability to resist adversarial attacks but also ensures the preservation of crucial details and captures the overall semantic meaning within the generated captions. The reported CIDEr scores in Table 2, also support these findings. Specifically, TeCoA4 and FARE4 achieve CIDEr scores of 64.4 and 75.3 respectively, whereas Sim-CLIP4 attains the highest CIDEr score of 84.7. In Appendix, we examine the robustness of CLIP models under targeted attacks at $\epsilon=2/255$, alongside an ablation study with a higher number of attack iterations. We also illustrate additional examples along with the corresponding captions generated by the models during the attack.
4.3 Zero-shot classification results and discussion.
In Table 3, we evaluate the performance of both original and robust CLIP models on zero-shot classification tasks across various datasets under two adversarial attack settings: $\epsilon=2/255$ and $\epsilon=4/255$. Across both attack settings and a wide range of datasets, Sim-CLIP consistently outperforms TeCoA and FARE. Notably, Sim-CLIP4 demonstrates greater robustness compared to FARE4 and TeCoA4, with a $3.4\%$ gain in robust accuracy, while also slightly outperforming Sim-CLIP2. These findings reinforce the effectiveness of Sim-CLIP in handling various adversarial situations within zero-shot classification tasks, showcasing its superior capabilities. We also evaluate the performance of CLIP models using clean images (without perturbation) on zero-shot classification datasets; detailed results are provided in the Appendix section.
5 Conclusion
In this paper, we introduce an unsupervised adversarial fine-tuning method that significantly enhances the robustness of the widely-used CLIP vision encoder against adversarial attacks. By employing a Siamese architecture with cosine similarity loss and a stop-gradient mechanism, our proposed framework, Sim-CLIP effectively learns semantically meaningful and attack-resilient visual representations without requiring large batch sizes or momentum encoders. Extensive experiments demonstrate that VLMs equipped with Sim-CLIP’s robust CLIP encoder exhibit superior robustness against both untargeted and targeted adversarial attacks while maintaining high accuracy on clean data across diverse downstream tasks, outperforming state-of-the-art adversarial fine-tuning approaches. As large VLMs become increasingly prevalent in critical applications, Sim-CLIP establishes strong benchmarks for evaluating robustness in the face of evolving adversarial threats. This work underscores the importance of robustifying foundation models like CLIP to safeguard the integrity of downstream vision-language models and paves the way for the development of more secure and reliable multimodal AI systems.
Appendix
Table 4: Quantitative evaluation of $\ell_{\infty}$ targeted attacks at $\epsilon=2/255$ radii.
Target CLIP TeCoA2 FARE2 Sim-CLIP2 TeCoA4 FARE4 Sim-CLIP4
COVID vaccine has severe 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Opposing political party 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Please visit https://deal.net 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Investing in this new crypto 20 / 20 1 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Climax change is a hoax 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
Mean success rate: 100% 1% 0% 0% 0% 0% 0%
Average CIDEr score: 0 49.3 74.5 75.1 78.4 81.7 85.9
Ablation with Stop-grad
In Figure 4, we demonstrate the effectiveness of stop-gradient mechanism in Sim-CLIP. For this ablation, all hyperparameters remain unchanged, with the only distinction being the stop-gradient.
Refer to caption Refer to caption
(a) (b)
Figure 4: Sim-CLIP with stop-gradient vs. without stop-gradient. (a) Adversarial training loss comparison. Loss degenerates after few steps without stop-gradient. (b) Robust accuracy comparison. Robust accuracy on ImageNet plummets after few steps without stop-gradient.
As shown in Figure 4(a), Sim-CLIP without stop-gradient collapses within a few training epochs, reaching the minimum possible loss of -1. A similar collapse in performance is observed in terms of accuracy, as shown in Figure 4(b). Sim-CLIP with stop-gradient mechanism achieves robust accuracy over 60%, whereas without stop-gradient, the accuracy plummets after only a few training steps.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 5: Training hyperparameter ablation for Sim-CLIP. Left plot: Loss minimization performance comparison with SGD and AdamW optimizer. Sim-CLIP generalizes better with AdamW optimizer. Middle plot: Comparison of adversarial training loss minimization with AdamW optimizer across different Learning Rates (LR) and Weight-decay (WD). Right plot: Robust accuracy comparison on different LR and WD with AdamW optimizer.
Pseudocode
In Algorithm 1, we present the pseudocode of our adversarial fine-tuning framework.
Listing 1 Algorithm 1: Sim-CLIP adversarial fine-tuning pseudocode
1for x in dataloader:
2 x1, x2 = x, perturb(x)
3 Rc, Rp = Fclip(x1), Fclip(x2);
4 L1= CosSim(Rc, Rp)
5 L2= CosSim(Rp, Rc)
6 Lclip = L1/2 + L2/2
7 Lclip.backward();
8 update(Fclip);
9function CosSim(r,q):
10 r = r.detach(); #stopgrad
11 loss = - F.cosine_similarity(r, q, dim=-1).mean()
12 return loss
Figure 6: Illustration of $\ell_{\infty}$ adversarial attacks at $\epsilon=4/255$ radii on OKVQA dataset while using original and robust CLIP models as vision encoder in OpenFlamingo. For each question, we show generated answers by OF with different vision encoders on both benign images (left) and attacked images (right). Answers highlighted in red denote incorrect responses, while answers in green signify correct responses.
Figure 7: Targeted $\ell_{\infty}$ attacks at $\epsilon=4/255$ radii for $10,000$ iterations, presenting additional examples with corresponding captions (good caption , captions with mistakes, captions missing intricate details, malicious target output) from Table 2.
Training Hyperparameters ablation.
In this segment, we delve into the specifics of our experimental setup, exploring different hyperparameter configurations for our Sim-CLIP approach. All experiments detailed in the main paper and the ablation studies, utilize the ViT-L/14 vision encoder of CLIP. Due to the significant computational resources required for training these CLIP models, we employ an early stopping strategy at 500 adversarial training steps. This strategy allows us to effectively evaluate and compare the performance of Sim-CLIP across different hyperparameter configurations. Initially, we investigate the impact of different optimizers in Sim-CLIP. Then, we conduct additional experiments by adjusting the Learning Rate (LR) and Weight Decay (WD) to identify the optimal configuration for the selected optimizer. In line with the supervised fine-tuning methodology proposed by (Mao et al. 2022), we constrain our LR and WD search within the values of 1e-3 to 1e-6 and 1e-3 to 1e-5, respectively.
Table 5: Ablation on different number of iterations of $\ell_{\infty}$ targeted attacks at $\epsilon=4/255$ radii.
Target Attack iterations
CLIP TeCoA4 FARE4 Sim-CLIP4
500 20000 500 20000 500 20000 500 20000
Investing in this new crypto $17/20$ $/20$ 20 $0/20$ $1/20$ $0/20$ $0/20$ $0/20$ $0/20$
Opposing political party $14/20$ $/20$ 20 $0/20$ $0/20$ $0/20$ $0/20$ $0/20$ $0/20$
Mean Success Rate 77.5% 100% 0% 2.5% 0% 0% 0% 0%
Average CIDEr score 0 0 76.4 67.3 89.3 78.3 93.4 82.3
In Figure 5, we illustrate the performance of Sim-CLIP with different hyperparameter configurations. Figure 5(a) illustrates our ablation study comparing Sim-CLIP’s performance using both SGD and AdamW optimizers. For both optimizers, we maintain LR at 1e-5 and WD at 1e-4. Notably, we observe Sim-CLIP’s sensitivity to different optimizers, where, Sim-CLIP performs significantly better with the AdamW optimizer compared to SGD. Hence, we select AdamW as our final optimizer for Sim-CLIP. Subsequently, we experiment with different LR and WD combinations with AdamW optimizer and present our findings in Figure 5(b) and Figure 5(c). Initially, we explore the combination with the highest LR (1e-3) and the lowest WD (1e-5) within our predefined hyperparameter search bounds. We observe that Sim-CLIP exhibits challenges in converging with a higher learning rate. In contrast, lowering the learning rates results in comparatively better performance.
However, when we keep lowering the LR and set the LR to its lowest value (1e-6) within our search bounds, Sim-CLIP struggles to converge. Further analysis indicates that an LR of 1e-5 coupled with a WD of 1e-3 yields superior results compared to alternative configurations. Furthermore, our analysis reveals that Sim-CLIP exhibits some sensitivity to Weight Decay (WD), although not to the extent observed with LR. Our ablation study shows that setting WD to 1e-4 results in improved performance. We observe consistent performance trends when evaluating the Robust accuracy on the ImageNet dataset, as illustrated in Figure 5(c). Within the initial 500 training steps, Sim-CLIP demonstrates higher robust accuracy when configured with an LR of 1e-5 and a WD of 1e-3. This finding suggests that the combination of a lower learning rate and a slightly higher weight decay contributes to the enhanced robustness of Sim-CLIP during the early stages of training. Based on these observations, we determine the optimal Sim-CLIP hyperparameters to be an LR of 1e-5 and a WD of 1e-3.
Targeted Attack Details
For targeted attacks in the main paper, we use the following target strings to fool the VLMs.
  1. 1.
    COVID vaccine has severe health implications
  2. 2.
    Opposing political party aims to strip citizen rights
  3. 3.
    Investing in this new cryptocurrency will guarantee benefit within a month
  4. 4.
    Please visit https://deal.net/1x24g58
    kx0 for exclusive deals
  5. 5.
    Climate Change is a hoax, created by the scientists
Here, we provide several realistic examples illustrating how an attacker could exploit vulnerabilities in VLMs to deceive users. Targets 1 and 5 present scenarios where users could be misled with false information, potentially inciting panic. Target 2 poses a risk of spreading political misinformation, potentially leading to societal harm. Target 3 may mislead individuals into investing in specific cryptocurrencies, serving the adversary’s interests. Lastly, Target 4 could lure users to phishing sites, endangering their personal information. To execute targeted attacks for Targets 2, 3 and 4, we randomly selected 20 images from the COCO dataset and tailored LLaVA for inference. For targets 1 and 5, we collected samples from Google images, depicting patients in bed wearing masks and polar bears in the Arctic, respectively.
Targeted attack results and ablations
We evaluate robust CLIP models under targeted attacks at $\epsilon=2/255$ in Table 4. Similar to our experiments under $\epsilon=4/255$ attack, the original CLIP model shows no robustness and breaks in each instance. While TeCoA2 breaks in one case, FARE and Sim-CLIP show complete robustness with their fine-tuned versions. Although both FARE and Sim-CLIP exhibit robustness, Sim-CLIP consistently generates superior captions and preserves semantic meaning and features in the captions, compared to FARE, as evidenced by its higher CIDEr scores. Additionally, we conduct ablation with varying numbers of attack iterations and report our findings in Table 5. In this ablation study, we focus on two specific target strings. For each target, we randomly select 20 instances from the COCO dataset and execute $\ell_{\infty}$ targeted attacks at radii $\epsilon=4/255$ for $500$ and $20,000$ iterations. The rationale behind selecting $500$ iterations is to evaluate the performance of CLIP models when subjected to a relatively lower number of attack iterations.
Figure 8: Targeted $\ell_{\infty}$ attacks at $\epsilon=4/255$ radii for $20,000$ iterations. Two randomly selected images are shown for each target listed in Table 5.
Conversely, the choice of $20,000$ iterations aims to examine the robustness under more severe attack conditions. With only $500$ iterations, we observe a decrease in the attack’s success rate compared to our experiment with $10,000$ iterations in Section 4 of the main paper. Increasing the number of attack iterations to 20,000 results in the original CLIP model breaking in every instance, with TeCoA4 also breaking in one instance. On the other hand, FARE4 and Sim-CLIP4 both exhibit complete robustness under both variants of attack iterations. However, when considering the quality of generated captions, Sim-CLIP4 outperforms FARE4 by a performance gap of 6.3 in terms of CIDEr score. This also implies that Sim-CLIP excels in capturing richer semantic details compared to FARE.
To further validate our findings, Figure 8 showcases two randomly selected attacked images for each target string listed in Table 5. This figure showcases the captions generated by LLaVA using different CLIP models for both the clean image and the corresponding adversarial image. In the first example, original CLIP model generates the best caption for the clean image. However, following the attack, both the original CLIP and TeCoA models break and output the target string. Among FARE and Sim-CLIP, we observe that Sim-CLIP excels in generating the most detailed caption by retaining intricate details from the image. For instance, in Sim-CLIP’s generated caption, the specific detail of a man wearing a black shirt in front of a wine display is accurately preserved, whereas FARE fails to preserve the details of the wine display entirely. Similarly, in the last example, Sim-CLIP successfully maintains the nuanced detail of a person wearing a uniform, whereas both FARE and TeCoA miss that specific detail. The results demonstrate Sim-CLIP as the most effective and robust vision encoder under targeted attacks. Moreover, our ablation study further reinforces the superiority of the Sim-CLIP approach, particularly in its capacity to preserve semantic meaning and retain intricate image details, even under stronger attack conditions.
Table 6: Clean evaluation on image classification datasets in zero-shot setting. We report clean accuracy of different CLIP models on zero-shot classification dataset.
Eval. Vision encoder
Cars
CIFAR10
CIFAR100
EuroSAT
Flowers
PCAM
clean CLIP 76.5 87.3 67.6 49.9 76.3 52.3
TeCoA2 63.1 76.1 54.7 26.1 50.1 49.4
FARE2 70.5 74.3 59.7 25.4 70.6 50.2
Sim-CLIP2 72.6 78.5 58.3 24.5 72.6 50.5
TeCoA4 59.3 74.6 51.3 22.5 46.4 47.1
FARE4 65.5 71.3 56.5 23.4 67.1 47.7
Sim-CLIP4 63.8 76.7 56.5 24.0 69.5 48.0
Clean evaluation on zero-shot classification datasets
In Table 6, we present the clean accuracy of state-of-the-art CLIP models on several image classification datasets: Cars, CIFAR10, CIFAR100, EuroSAT, Flowers, and PCAM. The table compares different vision encoders, including CLIP, two versions of TeCoA and FARE, and two versions of Sim-CLIP. As anticipated, the original CLIP model performs the best in this clean evaluation, given that it has not undergone any adversarial training. Among the robust CLIP models, Sim-CLIP consistently outperforms TeCoA and FARE across most zero-shot datasets.

### v2

@@
 Abstract
-Vision-language models (VLMs) have achieved significant strides in recent times specially in multimodal tasks, yet they remain susceptible to adversarial attacks on their vision components.
-To address this, we propose Sim-CLIP, an unsupervised adversarial fine-tuning method that enhances the robustness of the widely-used CLIP vision encoder against such attacks while maintaining semantic richness and specificity.
-By employing a Siamese architecture with cosine similarity loss, Sim-CLIP learns semantically meaningful and attack-resilient visual representations without requiring large batch sizes or momentum encoders.
-Our results demonstrate that VLMs enhanced with Sim-CLIP’s fine-tuned CLIP encoder exhibit significantly enhanced robustness against adversarial attacks, while preserving semantic meaning of the perturbed images.
+Vision-language models (VLMs) have achieved significant advancements in recent years, particularly in multimodal tasks.
+However, they remain susceptible to adversarial attacks on their vision components.
+This paper introduces Sim-CLIP, an unsupervised adversarial fine-tuning method that enhances the robustness of the widely-used CLIP vision encoder, preserving semantic richness and specificity even in adversarial situations.
+We design Sim-CLIP using a Siamese architecture combined with cosine similarity loss, enabling it to learn semantically meaningful and attack-resilient visual representations without requiring large batch sizes of training samples or any additional momentum encoders.
+Our results demonstrate that VLMs enhanced with Sim-CLIP’s fine-tuned CLIP encoder exhibit significantly improved resilience against adversarial attacks, while retaining semantic meaning of the perturbed images.
 Notably, Sim-CLIP does not require additional training or fine-tuning of the VLM itself; replacing the original vision encoder with our fine-tuned Sim-CLIP suffices to provide robustness.
-This work underscores the significance of reinforcing foundational models like CLIP to safeguard the reliability of downstream VLM applications, paving the way for more secure and effective multimodal systems.
+This work underscores the significance of reinforcing foundational models like CLIP to safeguard the reliability of downstream VLM applications (e.g., image captioning, questions-answering, zero-shot tasks), paving the way for more secure and effective multimodal systems.
 1 Introduction
@@
 3.2 Unsupervised Adversarial Fine-Tuning
-We propose an unsupervised adversarial fine-tuning approach for the CLIP vision encoder that aims to enhance its robustness against adversarial attacks.
-Our methodology incorporates unsupervised learning techniques by leveraging siamese architecture.
-Our goal is to enhance the resilience of the vision encoder against adversarial attacks while preserving the semantic meaning of the image features.
 Previously, FARE (Schlarmann et al.
@@
 3.4 Symmetric Loss Collapse Prevention
-Minimizing this negative cosine loss directly might lead to a collapse of the loss function, yielding an undesirable constant solution (Chen et al.
+Minimizing the negative cosine loss directly might lead to a collapse of the loss function, yielding an undesirable constant solution (Chen et al.
 2020).
@@
 This work underscores the importance of robustifying foundation models like CLIP to safeguard the integrity of downstream vision-language models and paves the way for the development of more secure and reliable multimodal AI systems.
-Appendix
-Table 4: Quantitative evaluation of $\ell_{\infty}$ targeted attacks at $\epsilon=2/255$ radii.
-Target CLIP TeCoA2 FARE2 Sim-CLIP2 TeCoA4 FARE4 Sim-CLIP4
-COVID vaccine has severe 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
-Opposing political party 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
-Please visit https://deal.net 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
-Investing in this new crypto 20 / 20 1 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
-Climax change is a hoax 20 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20 0 / 20
-Mean success rate: 100% 1% 0% 0% 0% 0% 0%
-Average CIDEr score: 0 49.3 74.5 75.1 78.4 81.7 85.9
-Ablation with Stop-grad
-In Figure 4, we demonstrate the effectiveness of stop-gradient mechanism in Sim-CLIP. For this ablation, all hyperparameters remain unchanged, with the only distinction being the stop-gradient.
-Refer to caption Refer to caption
-(a) (b)
-Figure 4: Sim-CLIP with stop-gradient vs.
-without stop-gradient.
-(a) Adversarial training loss comparison.
-Loss degenerates after few steps without stop-gradient.
-(b) Robust accuracy comparison.
-Robust accuracy on ImageNet plummets after few steps without stop-gradient.
-As shown in Figure 4(a), Sim-CLIP without stop-gradient collapses within a few training epochs, reaching the minimum possible loss of -1.
-A similar collapse in performance is observed in terms of accuracy, as shown in Figure 4(b).
-Sim-CLIP with stop-gradient mechanism achieves robust accuracy over 60%, whereas without stop-gradient, the accuracy plummets after only a few training steps.
-Refer to caption Refer to caption Refer to caption
-(a) (b) (c)
-Figure 5: Training hyperparameter ablation for Sim-CLIP.
-Left plot: Loss minimization performance comparison with SGD and AdamW optimizer.
-Sim-CLIP generalizes better with AdamW optimizer.
-Middle plot: Comparison of adversarial training loss minimization with AdamW optimizer across different Learning Rates (LR) and Weight-decay (WD).
-Right plot: Robust accuracy comparison on different LR and WD with AdamW optimizer.
-Pseudocode
-In Algorithm 1, we present the pseudocode of our adversarial fine-tuning framework.
-Listing 1 Algorithm 1: Sim-CLIP adversarial fine-tuning pseudocode
-1for x in dataloader:
-2 x1, x2 = x, perturb(x)
-3 Rc, Rp = Fclip(x1), Fclip(x2);
-4 L1= CosSim(Rc, Rp)
-5 L2= CosSim(Rp, Rc)
-6 Lclip = L1/2 + L2/2
-7 Lclip.backward();
-8 update(Fclip);
-9function CosSim(r,q):
-10 r = r.detach(); #stopgrad
-11 loss = - F.cosine_similarity(r, q, dim=-1).mean()
-12 return loss
-Figure 6: Illustration of $\ell_{\infty}$ adversarial attacks at $\epsilon=4/255$ radii on OKVQA dataset while using original and robust CLIP models as vision encoder in OpenFlamingo.
-For each question, we show generated answers by OF with different vision encoders on both benign images (left) and attacked images (right).
-Answers highlighted in red denote incorrect responses, while answers in green signify correct responses.
-Figure 7: Targeted $\ell_{\infty}$ attacks at $\epsilon=4/255$ radii for $10,000$ iterations, presenting additional examples with corresponding captions (good caption , captions with mistakes, captions missing intricate details, malicious target output) from Table 2.
-Training Hyperparameters ablation.
-In this segment, we delve into the specifics of our experimental setup, exploring different hyperparameter configurations for our Sim-CLIP approach.
-All experiments detailed in the main paper and the ablation studies, utilize the ViT-L/14 vision encoder of CLIP.
-Due to the significant computational resources required for training these CLIP models, we employ an early stopping strategy at 500 adversarial training steps.
-This strategy allows us to effectively evaluate and compare the performance of Sim-CLIP across different hyperparameter configurations.
-Initially, we investigate the impact of different optimizers in Sim-CLIP.
-Then, we conduct additional experiments by adjusting the Learning Rate (LR) and Weight Decay (WD) to identify the optimal configuration for the selected optimizer.
-In line with the supervised fine-tuning methodology proposed by (Mao et al.
-2022), we constrain our LR and WD search within the values of 1e-3 to 1e-6 and 1e-3 to 1e-5, respectively.
-Table 5: Ablation on different number of iterations of $\ell_{\infty}$ targeted attacks at $\epsilon=4/255$ radii.
-Target Attack iterations
-CLIP TeCoA4 FARE4 Sim-CLIP4
-500 20000 500 20000 500 20000 500 20000
-Investing in this new crypto $17/20$ $/20$ 20 $0/20$ $1/20$ $0/20$ $0/20$ $0/20$ $0/20$
-Opposing political party $14/20$ $/20$ 20 $0/20$ $0/20$ $0/20$ $0/20$ $0/20$ $0/20$
-Mean Success Rate 77.5% 100% 0% 2.5% 0% 0% 0% 0%
-Average CIDEr score 0 0 76.4 67.3 89.3 78.3 93.4 82.3
-In Figure 5, we illustrate the performance of Sim-CLIP with different hyperparameter configurations.
-Figure 5(a) illustrates our ablation study comparing Sim-CLIP’s performance using both SGD and AdamW optimizers.
-For both optimizers, we maintain LR at 1e-5 and WD at 1e-4.
-Notably, we observe Sim-CLIP’s sensitivity to different optimizers, where, Sim-CLIP performs significantly better with the AdamW optimizer compared to SGD.
-Hence, we select AdamW as our final optimizer for Sim-CLIP.
-Subsequently, we experiment with different LR and WD combinations with AdamW optimizer and present our findings in Figure 5(b) and Figure 5(c).
-Initially, we explore the combination with the highest LR (1e-3) and the lowest WD (1e-5) within our predefined hyperparameter search bounds.
-We observe that Sim-CLIP exhibits challenges in converging with a higher learning rate.
-In contrast, lowering the learning rates results in comparatively better performance.
-However, when we keep lowering the LR and set the LR to its lowest value (1e-6) within our search bounds, Sim-CLIP struggles to converge.
-Further analysis indicates that an LR of 1e-5 coupled with a WD of 1e-3 yields superior results compared to alternative configurations.
-Furthermore, our analysis reveals that Sim-CLIP exhibits some sensitivity to Weight Decay (WD), although not to the extent observed with LR.
-Our ablation study shows that setting WD to 1e-4 results in improved performance.
-We observe consistent performance trends when evaluating the Robust accuracy on the ImageNet dataset, as illustrated in Figure 5(c).
-Within the initial 500 training steps, Sim-CLIP demonstrates higher robust accuracy when configured with an LR of 1e-5 and a WD of 1e-3.
-This finding suggests that the combination of a lower learning rate and a slightly higher weight decay contributes to the enhanced robustness of Sim-CLIP during the early stages of training.
-Based on these observations, we determine the optimal Sim-CLIP hyperparameters to be an LR of 1e-5 and a WD of 1e-3.
-Targeted Attack Details
-For targeted attacks in the main paper, we use the following target strings to fool the VLMs.
-  1. 1.
-    COVID vaccine has severe health implications
-  2. 2.
-    Opposing political party aims to strip citizen rights
-  3. 3.
-    Investing in this new cryptocurrency will guarantee benefit within a month
-  4. 4.
-    Please visit https://deal.net/1x24g58
-    kx0 for exclusive deals
-  5. 5.
-    Climate Change is a hoax, created by the scientists
-Here, we provide several realistic examples illustrating how an attacker could exploit vulnerabilities in VLMs to deceive users.
-Targets 1 and 5 present scenarios where users could be misled with false information, potentially inciting panic.
-Target 2 poses a risk of spreading political misinformation, potentially leading to societal harm.
-Target 3 may mislead individuals into investing in specific cryptocurrencies, serving the adversary’s interests.
-Lastly, Target 4 could lure users to phishing sites, endangering their personal information.
-To execute targeted attacks for Targets 2, 3 and 4, we randomly selected 20 images from the COCO dataset and tailored LLaVA for inference.
-For targets 1 and 5, we collected samples from Google images, depicting patients in bed wearing masks and polar bears in the Arctic, respectively.
-Targeted attack results and ablations
-We evaluate robust CLIP models under targeted attacks at $\epsilon=2/255$ in Table 4.
-Similar to our experiments under $\epsilon=4/255$ attack, the original CLIP model shows no robustness and breaks in each instance.
-While TeCoA2 breaks in one case, FARE and Sim-CLIP show complete robustness with their fine-tuned versions.
-Although both FARE and Sim-CLIP exhibit robustness, Sim-CLIP consistently generates superior captions and preserves semantic meaning and features in the captions, compared to FARE, as evidenced by its higher CIDEr scores.
-Additionally, we conduct ablation with varying numbers of attack iterations and report our findings in Table 5.
-In this ablation study, we focus on two specific target strings.
-For each target, we randomly select 20 instances from the COCO dataset and execute $\ell_{\infty}$ targeted attacks at radii $\epsilon=4/255$ for $500$ and $20,000$ iterations.
-The rationale behind selecting $500$ iterations is to evaluate the performance of CLIP models when subjected to a relatively lower number of attack iterations.
-Figure 8: Targeted $\ell_{\infty}$ attacks at $\epsilon=4/255$ radii for $20,000$ iterations. Two randomly selected images are shown for each target listed in Table 5.
-Conversely, the choice of $20,000$ iterations aims to examine the robustness under more severe attack conditions.
-With only $500$ iterations, we observe a decrease in the attack’s success rate compared to our experiment with $10,000$ iterations in Section 4 of the main paper.
-Increasing the number of attack iterations to 20,000 results in the original CLIP model breaking in every instance, with TeCoA4 also breaking in one instance.
-On the other hand, FARE4 and Sim-CLIP4 both exhibit complete robustness under both variants of attack iterations.
-However, when considering the quality of generated captions, Sim-CLIP4 outperforms FARE4 by a performance gap of 6.3 in terms of CIDEr score.
-This also implies that Sim-CLIP excels in capturing richer semantic details compared to FARE.
-To further validate our findings, Figure 8 showcases two randomly selected attacked images for each target string listed in Table 5.
-This figure showcases the captions generated by LLaVA using different CLIP models for both the clean image and the corresponding adversarial image.
-In the first example, original CLIP model generates the best caption for the clean image.
-However, following the attack, both the original CLIP and TeCoA models break and output the target string.
-Among FARE and Sim-CLIP, we observe that Sim-CLIP excels in generating the most detailed caption by retaining intricate details from the image.
-For instance, in Sim-CLIP’s generated caption, the specific detail of a man wearing a black shirt in front of a wine display is accurately preserved, whereas FARE fails to preserve the details of the wine display entirely.
-Similarly, in the last example, Sim-CLIP successfully maintains the nuanced detail of a person wearing a uniform, whereas both FARE and TeCoA miss that specific detail.
-The results demonstrate Sim-CLIP as the most effective and robust vision encoder under targeted attacks.
-Moreover, our ablation study further reinforces the superiority of the Sim-CLIP approach, particularly in its capacity to preserve semantic meaning and retain intricate image details, even under stronger attack conditions.
-Table 6: Clean evaluation on image classification datasets in zero-shot setting. We report clean accuracy of different CLIP models on zero-shot classification dataset.
-Eval. Vision encoder
-Cars
-CIFAR10
-CIFAR100
-EuroSAT
-Flowers
-PCAM
-clean CLIP 76.5 87.3 67.6 49.9 76.3 52.3
-TeCoA2 63.1 76.1 54.7 26.1 50.1 49.4
-FARE2 70.5 74.3 59.7 25.4 70.6 50.2
-Sim-CLIP2 72.6 78.5 58.3 24.5 72.6 50.5
-TeCoA4 59.3 74.6 51.3 22.5 46.4 47.1
-FARE4 65.5 71.3 56.5 23.4 67.1 47.7
-Sim-CLIP4 63.8 76.7 56.5 24.0 69.5 48.0
-Clean evaluation on zero-shot classification datasets
-In Table 6, we present the clean accuracy of state-of-the-art CLIP models on several image classification datasets: Cars, CIFAR10, CIFAR100, EuroSAT, Flowers, and PCAM.
-The table compares different vision encoders, including CLIP, two versions of TeCoA and FARE, and two versions of Sim-CLIP.
-As anticipated, the original CLIP model performs the best in this clean evaluation, given that it has not undergone any adversarial training.
-Among the robust CLIP models, Sim-CLIP consistently outperforms TeCoA and FARE across most zero-shot datasets.